# {textblock*}

### v1

@@ 3.3.3 EDA Script Generation and Architecture Design
Architectural Design: GPT4AIGChip [181] zeroes in on AI accelerator design, leveraging LLMs to generate C code for high-level synthesis. Similarly, Yan et al.[182] examine the use of LLMs in optimizing compute-in-memory (CiM) DNN accelerators, showcasing the model’s potential in enhancing computational efficiency. Further extending the scope, Liang et al.[183] delve into quantum architecture design, exploring the frontiers of quantum computing. SpecLLM [184] contributes to this growing body of work by providing a dataset of architecture specifications at various abstraction levels, investigating LLMs’ capabilities in both generating and reviewing these specifications.
@@ 4 Large Circuit Models: A New Horizon
Venturing into the domain of large circuit models (refer to Figure 1) marks a bold departure from them, moving towards a more integrated and AI-native design process. The term ‘large’ in LCMs signifies both the substantial model size and the vast array of circuit data collected from various EDA stages for circuit pre-training. Such a foundational model concept promises a unified framework that transcends task-oriented limitations, ensuring that LCMs are robust, versatile, and capable of handling the diverse tasks of modern circuit design with limited fine-tuning.
@@ 4.2 Overview of LCMs
Addressing the complexities of this multimodal fusion, the perspective paper outlines three primary alignment challenges:
@@ 4.3 Promises and Potentials
5 Unimodal Circuit Representation
Learning
@@ 5.1.1 Representation Learning for Architecture Design
The performance and power consumption of architectures exhibit an intrinsic dependence on specific application contexts. In pursuit of optimizing the trade-off among performance, power efficiency, and silicon area for targeted applications, architectural designers traditionally employ detailed simulation tools complemented by extensive domain-specific expertise. This conventional methodology, while comprehensive, tends to be both time-intensive and prone to human errors. The advent of LCM presents a novel paradigm, facilitating rapid exploration of architectural design spaces by leveraging insights into the nuanced interactions between application workloads and architectural configurations. Thus, it is imperative for LCM to encapsulate application workload representations adaptable to various architectural designs.
@@ 5.1.2 Representation Learning for HLS/RTL
To incorporate deep learning in understanding and optimizing HLS/RTL representations, we can explore two innovative methodologies. The first method interprets code as a series of tokens, analogous to words in natural language, making it possible to apply NLP techniques to HLS/RTL codes. A particularly effective strategy in this domain is masked language modeling (MLM), where certain tokens are obscured during model training, prompting a Transformer-based encoder (such as BERT) to infer the missing tokens. This self-supervised learning approach yields representations rich in the semantic essence of the hardware design, capturing the functional nuances at both the HLS and RTL levels.
Alternatively, representing HLS/RTL designs as control data flow graphs (CDFGs) offers a graphical perspective, mapping out the control and data dependencies within the design. Here, advanced GNNs come into play, learning from the complex web of interactions and dependencies depicted in the CDFGs. This method allows for the extraction of comprehensive representations that embody the intricate structure and operational logic of the design, providing a solid foundation for subsequent optimization and synthesis tasks.
The token view is more aligned with the high-level specifications and contains more syntax information. With the application of language models that excel in capturing global relationships, we can get representations that encompass the overall behavior and functionality of the design. Besides, the learned representations will benefit the generalizability and scalability of the attention-based models. On the other hand, the graph view is more aligned with the lower-level gate-level representations and contains more structural and semantic information. Compared to language models, GNNs focus more on extracting local information.
@@ 5.1.3 Representation Learning for Circuit Netlist
Figure 7: DeepGate2: Structural and Functional Disentangled Netlist Representation Learning
@@
While the primary focus of logic synthesis has been on optimizing combinational logic, the sequential behavior of circuits is also a critical facet to represent. DeepSeq [105] expands upon the DeepGate technique by elucidating the temporal correlations within sequential netlists. This advancement is facilitated by leveraging both transition and logic-1 probabilities for supervision across each logic gate and memory element, where transition probabilities unveil insights into the circuit’s state transition behaviors and logic-1 probabilities illuminate functional and topological characteristics. Such a nuanced approach allows DeepSeq to adeptly encode the complex dynamics and behaviors of sequential circuits, proving instrumental for downstream applications such as netlist-level power estimation and reliability analysis
@@ 5.2 Representation Learning for Back-End Design
The meticulous development of unimodal representations across each design stage knits a rich tapestry of circuit knowledge. Existing studies have explored unimodal learning for prediction of various factors such as routability, IR drop, and lithography hotspots [113, 224, 134]. Although back-end design consists of many design stages with different levels of geometric abstraction, as shown in Figure 3, existing studies mostly focus on individual stages. There are a few key problems yet to be solved before making back-end representation learning practical in real design applications. For easier understanding, we interpret the layout representation learning task by comparing with computer vision tasks on images.
@@ 6.1 Implementing Multimodal Circuit Alignment
Central to the concept of multimodal circuit learning is the understanding that all design stages, although distinct in form, share a common functional objective. By applying sophisticated feature extraction and alignment techniques, it becomes possible to overcome the semantic disconnects that typically arise in representation learning. This ensures that the original design intent is not only preserved but also accentuated throughout the entire design lifecycle. The adoption of machine learning models, particularly those leveraging attention mechanisms and joint embedding spaces, promises to lead the charge towards a more integrated and holistic approach to circuit design.
@@
A potential solution to achieve this alignment involves the use of masked modeling across different modalities. This technique, inspired by successful applications in natural language processing, involves selectively hiding parts of the input data across modalities and then training the model to predict these masked portions. By applying this method across circuit design representations—ranging from natural language specifications, high-level algorithms, and RTL implementations to detailed physical layouts can learn a joint representation that captures the essence of the design process at various abstraction levels. This joint representation is crucial for the model to understand the transition from high-level specifications to detailed implementations, enabling it to navigate the complexities of circuit design with greater precision and efficiency.
@@
Addressing the variability in how a high-level design can be mapped to multiple lower-level implementations, each with PPA characteristics, poses a significant challenge. To tackle this, models need to be equipped with the ability to recognize and evaluate the trade-offs associated with different design choices. Integrating reinforcement learning techniques with the multimodal learning framework can provide a solution. By setting the optimization of PPA metrics as the reward function, the model can learn to navigate the space of possible implementations, identifying solutions that best meet the specified criteria. Furthermore, incorporating attention mechanisms can enhance the model’s ability to focus on relevant features across modalities, thereby improving its capacity to predict implementations that not only meet functional requirements but also optimize for PPA objectives. Through these methods, the implementation of multimodal alignment in circuit design can become not just a theoretical concept but a practical tool for advancing the field.
@@
Considering the vast differences between design modalities, aligning them in a single step is a formidable challenge. To address this, we propose a phased approach to multimodal alignment, partitioning the process into three distinct phases: “Spec-HLS-RTL Representation Alignment,” “RTL-Netlist Representation Alignment,” and “Netlist-Layout Representation Alignment.” This strategy employs RTL and circuit netlists as intermediaries, facilitating a more manageable and focused alignment process. By breaking down the alignment into these stages, we can concentrate on specific transitions within the design flow, allowing for a more tailored application of machine learning techniques to each phase. This phased approach not only makes the task of alignment more feasible but also ensures that each stage of the design process is optimally aligned, leading to more coherent and efficient design outcomes. Through careful implementation of this strategy, we aim to bridge the gap between the various design modalities, ultimately fostering a more integrated and seamless circuit design environment.
@@ 7 Pioneering LCM Applications
TABLE I: Solving Time Comparison between Ours and [226] on LEC Cases
@@ 7.1 Circuit Learning for SAT
The Boolean Satisfiability (SAT) problem identifies if there exists at least one assignment that makes a given Boolean formula to be True. SAT problem acts as a fundamental problem in many areas, especially in the EDA fields, such as logic equivalence checking, model checking and testing. Over the past few decades, the SAT community has advocated adopting the conjunctive normal form (CNF) as the de facto standard format for problem instances and developed numerous advanced CNF-based SAT solvers [227, 228]. However, the efficacy of CNF-based solvers recently encounter bottlenecks in solving hard SAT problems, prompting past research to explore circuit-based solvers or strategies as a potential breakthrough. In this section, we aim to demonstrate the impact of the large circuit model on SAT solving.
@@ 8.3 Large Circuit Models for Analog Circuits
TAG [192] represents an early effort to develop a circuit representation model for analog EDA. It introduces a netlist embedding mechanism and a ”pretrain-then-finetune” strategy to apply embedding vectors across various applications. However, it lacks a unified, aligned representation across all design stages, with its effectiveness constrained by the initial pretraining target, layout distance. Its potential applications are somewhat limited compared to a comprehensive LCM for analog circuits.
Figure 8 presents our vision for future analog large circuit models. These models are inherently multimodal, capable of processing various data structures from different design stages. Text and graph structures can represent a netlist, while images may be used for layout designs. An Analog LCM converts these inputs into vectors, mapping the designs to a unified embedding space. The generated circuit embedding vectors can then support various downstream tasks across different circuit types (such as amplifiers, PLLs, and ADCs), catering to a range of applications from topology design to routing.
@@ 9.1 Data Issues
Data scarcity stands out as a critical hurdle, given the dependency of LCMs on extensive, high-quality datasets for training. The realm of circuit design, particularly at the granularity required for effective LCM operation, suffers from a lack of publicly available data, posing risks of overfitting and undermining the models’ generalization capabilities. Tackling this issue head-on, innovative data augmentation techniques emerge as a key solution. For instance, equivalent circuits can be generated through systematic circuit transformations, effectively expanding the dataset without the need for additional real-world data. This approach not only enhances the diversity of training material but also deepens the model’s understanding of circuit variability and design principles.
On the synthetic data front, leveraging LLMs to generate realistic RTL code presents an exciting opportunity. This strategy involves using LLMs’ advanced generative capabilities to create new RTL designs, which can then serve both as additional training data and as benchmarks for further refining the RTL generative models themselves. This creates a self-reinforcing loop where LCMs continually improve through iterative training on both real and synthetically generated data. Such a mechanism not only addresses the issue of data scarcity but also contributes to the evolution of more sophisticated and capable generative models, marking a significant leap towards fully realizing the transformative potential of LCMs in the EDA landscape.
Furthermore, the development of community-driven platforms for data sharing and collaboration could significantly alleviate the scarcity issue. By fostering an ecosystem where academia and industry share circuit data and design challenges, the field can collectively advance the state of LCM research, ensuring a diverse and comprehensive dataset that mirrors the multifaceted nature of electronic design automation.
@@ 9.3 New Opportunities
Finally, the synergy between large language models and LCMs presents a particularly promising area of exploration. LLMs, with their advanced natural language processing capabilities, can serve as intuitive, conversational interfaces for designers, translating high-level design specifications into actionable insights and suggestions; While the LCMs, with their deep understanding of circuitry and design principles, can analyze and optimize the granular details of the netlist, ensuring that the final design aligns with the desired performance, power, and area constraint. This collaborative interaction between LLMs and LCMs allows for a seamless transition from abstract design concepts to concrete, optimized circuit representations. Bridging the gap between high-level design intent and detailed technical execution, this synergy enables a more holistic and integrated approach to circuit design.

### v2

@@ 3.3.3 EDA Script Generation and Architecture Design
Architectural Design: GPT4AIGChip [184] leverages LLMs to generate C code for AI accelerator high-level synthesis. Similarly, Yan et al.[185] examine the use of LLMs in optimizing compute-in-memory (CiM) DNN accelerators, showcasing the model’s potential in enhancing computational efficiency. Further extending the scope, Liang et al.[186] delve into quantum architecture design, exploring the frontiers of quantum computing. SpecLLM [187] contributes to this growing body of work by providing a dataset of architecture specifications at various abstraction levels, investigating LLMs’ capabilities in both generating and reviewing these specifications.
@@ 4 Large Circuit Models: A New Horizon
Venturing into the domain of large circuit models (refer to Fig. 1) marks a bold departure from the previous AI4EDA solutions, moving towards a more integrated and AI-native design process. The term ‘large’ in LCMs signifies both the substantial model size and the vast array of circuit data collected from various EDA stages for circuit pre-training. Such a foundational model concept promises a unified framework that transcends task-oriented limitations, ensuring that LCMs are robust, versatile, and capable of handling the diverse tasks of modern circuit design with limited fine-tuning.
@@ 4.2 Overview of LCMs
Since the specifications, RTL codes, netlists and layout designs are representative formats in front-end and back-end flows, the perspective paper outlines three primary alignment challenges:
@@ 4.3 Opportunities and Potentials
5 Unimodal Circuit Representation Learning
@@ 5.1.1 Representation Learning for Architecture Design
The performance and power consumption of architectures exhibit an intrinsic dependence on specific application contexts. In pursuit of optimizing the trade-off among PPA for targeted applications, architectural designers traditionally employ detailed simulation tools complemented by extensive domain-specific expertise. This conventional methodology, while comprehensive, tends to be both time-intensive and prone to human errors. The advent of LCM presents a novel paradigm, facilitating rapid exploration of architectural design spaces by leveraging insights into the nuanced interactions between application workloads and architectural configurations. Thus, it is imperative for LCM to encapsulate application workload representations adaptable to various architectural designs.
@@ 5.1.2 Representation Learning for HLS/RTL
To incorporate deep learning in understanding and optimizing HLS/RTL representations, we can explore two innovative methodologies. One method interprets code as a series of tokens, analogous to words in natural language, making it possible to apply NLP techniques to HLS/RTL codes. A particularly effective strategy in this domain is masked language modeling (MLM), where certain tokens are obscured during model training, prompting a Transformer-based encoder (such as BERT) to infer the missing tokens. This self-supervised learning approach yields representations rich in the semantic essence of the hardware design, capturing the functional nuances at both the HLS and RTL levels. Another method may represent HLS/RTL designs as control data flow graphs (CDFGs) offers a graphical perspective, mapping out the control and data dependencies within the design. Here, advanced GNNs come into play, learning from the complex web of interactions and dependencies depicted in the CDFGs. This method allows for the extraction of comprehensive representations that embody the intricate structure and operational logic of the design, providing a solid foundation for subsequent optimization and synthesis tasks.
The former token view is more aligned with the high-level specifications and contains more syntax information. With the application of language models that excel in capturing global relationships, we can get representations that encompass the overall behavior and functionality of the design. Besides, the learned representations will benefit the generalizability and scalability of the attention-based models. On the other hand, the graph view is more aligned with the lower-level gate-level representations and contains more structural and semantic information. Compared to language models, GNNs focus more on extracting local information.
@@ 5.1.3 Representation Learning for Circuit Netlist
Figure 6: DeepGate2: structural and functional disentangled netlist representation learning.
@@
While the primary focus of logic synthesis has been on optimizing combinational logic, the sequential behavior of circuits is also a critical facet to represent. DeepSeq [106] expands upon the DeepGate technique by elucidating the temporal correlations within sequential netlists. This advancement is facilitated by leveraging both transition and logic-1 probabilities for supervision across each logic gate and memory element, where transition probabilities unveil insights into the circuit’s state transition behaviors and logic-1 probabilities illuminate functional and topological characteristics. Such a nuanced approach allows DeepSeq to adeptly encode the complex dynamics and behaviors of sequential circuits, proving instrumental for downstream applications such as netlist-level power estimation and reliability analysis.
@@ 5.2 Representation Learning for Back-End Design
The meticulous development of unimodal representations across each design stage knits a rich tapestry of circuit knowledge. Existing studies have explored unimodal learning for prediction of various factors such as routability, IR drop, and lithography hotspots [114, 227, 135]. Although back-end design consists of many design stages with different levels of geometric abstraction, as shown in Fig. 2 b), existing studies mostly focus on individual stages. There are a few key problems yet to be solved before making back-end representation learning practical in real design applications. For easier understanding, we interpret the layout representation learning task by comparing with computer vision tasks on images.
@@ 6.1 Implementing Multimodal Circuit Alignment
Central to the concept of multimodal circuit learning is the understanding that all design stages, although distinct in form, share a common functional objective. By applying sophisticated feature extraction and alignment techniques, it becomes possible to overcome the semantic disconnects that typically arise in representation learning. This ensures that the original design intent is not only preserved but also accentuated throughout the entire design lifecycle. The adoption of machine learning models, particularly those leveraging scalable self-attention mechanisms and joint embedding spaces, promises to lead the charge towards a more integrated and holistic approach to circuit design.
@@
A potential solution to achieve this alignment involves the use of masked modeling across different modalities. This technique, inspired by successful applications [228] in natural language processing, involves selectively hiding parts of the input data across modalities and then training the model to predict these masked portions. By applying this method across circuit design representations—ranging from natural language specifications, high-level algorithms, and RTL implementations to detailed physical layouts can learn a joint representation that captures the essence of the design process at various abstraction levels. This joint representation is crucial for the model to understand the transition from high-level specifications to detailed implementations, enabling it to navigate the complexities of circuit design with greater precision and efficiency.
@@
However, addressing the variability in how a high-level design can be mapped to multiple lower-level implementations, each with PPA characteristics, poses a significant challenge. To tackle this, models need to be equipped with the ability to recognize and evaluate the trade-offs associated with different design choices. Integrating reinforcement learning techniques with the multimodal learning framework can provide a solution. By setting the optimization of PPA metrics as the reward function, the model can learn to navigate the space of possible implementations, identifying solutions that best meet the specified criteria. Furthermore, incorporating attention mechanisms can enhance the model’s ability to focus on relevant features across modalities, thereby improving its capacity to predict implementations that not only meet functional requirements but also optimize for PPA objectives. Through these methods, the implementation of multimodal alignment in circuit design can become not just a theoretical concept but a practical tool for advancing the field.
@@
Considering the vast differences between design modalities, aligning them in a single step is a formidable challenge. To address this, we propose a phased approach to multimodal alignment, partitioning the process into three distinct phases: “Spec-HLS-RTL Representation Alignment,” “RTL-Netlist Representation Alignment,” and “Netlist-Layout Representation Alignment.” Since the RTL design contains high-level semantics and netlist is more suitable for aligning with the following backend designs, this strategy employs these two designs as intermediaries, facilitating a more manageable and focused alignment process. By breaking down the alignment into these stages, we can concentrate on specific transitions within the design flow, allowing for a more tailored application of machine learning techniques to each phase. This phased approach not only makes the task of alignment more feasible but also ensures that each stage of the design process is optimally aligned, leading to more coherent and efficient design outcomes. Through careful implementation of this strategy, we aim to bridge the gap between the various design modalities, ultimately fostering a more integrated and seamless circuit design environment.
@@ 7 Pioneering LCM Applications
TABLE I: Solving time comparison between Ours and [230] on LEC cases.
@@ 7.1 Circuit Learning for SAT
The Boolean Satisfiability (SAT) problem identifies if there exists at least one assignment that makes a given Boolean formula to be True. SAT problem acts as a fundamental problem in many areas, especially in the EDA fields, such as logic equivalence checking, model checking, and testing. Over the past few decades, the SAT community has advocated adopting the conjunctive normal form (CNF) as the de facto standard format for problem instances and developed numerous advanced CNF-based SAT solvers [231, 232]. However, the efficacy of CNF-based solvers recently encounter bottlenecks in solving hard SAT problems, prompting past research to explore circuit-based solvers or strategies as a potential breakthrough. In this section, we aim to demonstrate the impact of the large circuit model on SAT solving.
@@ 8.3 Large Circuit Models for Analog Circuits
TAG [195] represents an early effort to develop a circuit representation model for analog EDA. It introduces a netlist embedding mechanism and a “pretrain-then-finetune” strategy to apply embedding vectors across various applications. However, it lacks a unified, aligned representation across all design stages, with its effectiveness constrained by the initial pretraining target, layout distance. Its potential applications are somewhat limited compared to a comprehensive LCM for analog circuits.
Fig. 7 presents our vision for future analog large circuit models. These models are inherently multimodal, capable of processing various data structures from different design stages. Text and graph structures can represent a netlist, while images may be used for layout designs. An Analog LCM converts these inputs into vectors, mapping the designs to a unified embedding space. The generated circuit embedding vectors can then support various downstream tasks across different circuit types (such as amplifiers, PLLs, and ADCs), catering to a range of applications from topology design to routing.
@@ 9.1 Data Issues
Data scarcity stands out as a critical hurdle, given the dependency of LCMs on extensive, high-quality datasets for training. The realm of circuit design, particularly at the granularity required for effective LCM operation, suffers from a lack of publicly available data, posing risks of overfitting and undermining the models’ generalization capabilities. Tackling this issue head-on, we introduce three possible solutions.
First, innovative data augmentation techniques emerge as a key solution. For instance, equivalent circuits can be generated through systematic circuit transformations, effectively expanding the dataset without the need for additional real-world data. This approach not only enhances the diversity of training material but also deepens the model’s understanding of circuit variability and design principles.
Second, on the synthetic data front, leveraging LLMs to generate realistic RTL code presents an exciting opportunity. This strategy involves using LLMs’ advanced generative capabilities to create new RTL designs, which can then serve both as additional training data and as benchmarks for further refining the RTL generative models themselves. This creates a self-reinforcing loop where LCMs continually improve through iterative training on both real and synthetically generated data. Such a mechanism not only addresses the issue of data scarcity but also contributes to the evolution of more sophisticated and capable generative models, marking a significant leap towards fully realizing the transformative potential of LCMs in the EDA landscape.
Third, the development of community-driven platforms for data sharing and collaboration could significantly alleviate the scarcity issue. By fostering an ecosystem where academia and industry share circuit data and design challenges, the field can collectively advance the state of LCM research, ensuring a diverse and comprehensive dataset that mirrors the multifaceted nature of electronic design automation.
@@ 9.3 New Opportunities
Finally, the synergy between large language models and LCMs presents a particularly promising area of exploration. LLMs, with their advanced natural language processing capabilities, can serve as intuitive, conversational interfaces for designers, translating high-level design specifications into actionable insights and suggestions; While the LCMs, with their deep understanding of circuitry and design principles, can analyze and optimize the granular details of the netlist, ensuring that the final design aligns with the desired performance, power, and area constraints. This collaborative interaction between LLMs and LCMs allows for a seamless transition from abstract design concepts to concrete, optimized circuit representations. Bridging the gap between high-level design intent and detailed technical execution, this synergy enables a more holistic and integrated approach to circuit design.